OpenStudy (anonymous):

solve each equation? a)5^3^x = 5^15 b)6^3^+^7^c = 6^2^c^-^3 c)2^5^w = 16

OpenStudy (cwrw238):

a) 3^x = 15 x log 3 = log 15 x = log 15 / log 3 = 2.46

OpenStudy (experimentx):

(5^3)^x = 5^15 or 5^(3^x) = 5^15 ??

OpenStudy (anonymous):

2^5^w = 16 2^5^w = 2^4 5^w=4 wln5=ln4 w=(ln4)/(ln5)

OpenStudy (anonymous):

\[\LARGE (5^3)^x=5^{15} \longrightarrow 5^{3x}=5^{3\cdot5} \longrightarrow x=??\]